What is the new boiling point of a solution of sugar water with a concentration of 1.25 m?


1
Expert's answer
2021-03-29T06:02:10-0400

Molar Mass of water = 18.01528

= 342 +18.01528 = 360.01528

343×1.5 = 513.5

= 513.5 - 360.01= 153.5

153/18.01528 = 8.5 +1 = 9.5 moles of H2O

Molar Mass of Sucrose = 342.3

= ∆Tf = Kf× m

Molality = 9.5/18.01528 = 0.527 mol/Kg

H2O has H+ and (OH)- ions

i = 2

Boiling point of water = 100.0°C

KbC(°C/m) = 0.52

= 0.52 ×2× 0.527 = 0.54808

= 0.54808 +100

= 100.55°C